1. The effectiveness of a blood-pressure drug is being investigated. An experimenter finds that, on average, the reduction in systolic blood pressure is 69.6 for a sample of size 27 and standard deviation 11.5. Estimate how much the drug will lower a typical patient's systolic blood pressure (using a 80% confidence level). Assume the data is from a normally distributed population.

Enter your answer as a tri-linear inequality accurate to three decimal places.

_____ < μ < _____

2. Express the confidence interval (25.5%,37.5%)(25.5%,37.5%) in the form of ˆp±Ep^±E.

______% ± ______%
Answer #1

Solution :

1)

t /2,df = 1.315

Margin of error = E = t/2,df * (s /n)

= 1.315 * (11.5 / 27)

Margin of error = E = 2.910

The 80% confidence interval estimate of the population mean is,

- E < < + E

69.6 - 2.910 < < 69.6 + 2.910

66.690 < < 72.510

2)

Point estimate = = (25.5% + 37.5%) / 2 = 31.5%

Margin of error = E = Upper confidence interval - = 37.5% - 31.5% = 6%

In the form of :

- E < p < + E is,

31.5% 6%